Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Tough Probability

Expert replies
by late4thing » Wed Oct 05, 2016 5:18 pm
Suppose Mr. Smith goes to Vegas. He knows from previous experiences that the probability he will come out ahead after playing black jack is .3. The probability he will be ahead after playing the roulette wheel is .4. He also knows that his chances of coming out ahead in one game is independent of the other. If he plays both games, find the probability he will end up being ahead in at least one.
Join the discussion
Source: — Problem Solving |

by crackverbal » Thu Oct 06, 2016 1:15 am
Hi late4thing,

Always remember that P(atleast one) = 1 - P(none)

Now the probability of Mr. Smith being ahead in blackjack is P(B) = 0.3. The probability of Mr. Smith not being ahead in blackjack will be P(not B) = 0.7.

The probability of Mr. Smith being ahead in roulette is P(R) = 0.4. The probability of Mr. Smith not being ahead in roulette will be P(not R) = 0.6.

The questions asks us to find the probability of Mr. Smith being ahead in at least one out of blackjack and roulette.

P(being ahead in atleast one) = 1 - P(being ahead in none)
= 1 - (0.7 * 0.6)
= 1 - 0.42
= 0.58

Hope this helps!

CrackVerbal Academics Team
Join Free 4 part MBA Through GMAT Video Training Series here -
https://gmat.crackverbal.com/mba-throug ... video-2018

Enroll for our GMAT Trial Course here -
https://gmatonline.crackverbal.com/

For more info on GMAT and MBA, follow us on @AskCrackVerbal
Join the discussion

by Brent@GMATPrepNow » Thu Oct 06, 2016 6:45 am
late4thing wrote:Suppose Mr. Smith goes to Vegas. He knows from previous experiences that the probability he will come out ahead after playing blackjack is .3. The probability he will be ahead after playing the roulette wheel is .4. He also knows that his chances of coming out ahead in one game is independent of the other. If he plays both games, find the probability he will end up being ahead in at least one.
crackverbal's approach is the same as the one I'd use. However, we can also solve the question this way:

P(ahead in blackjack) = 0.3
P(NOT ahead in blackjack) = 0.7
P(ahead in roulette) = 0.4
P(NOT ahead in roulette) = 0.6

P(ahead in at least one) = P(ahead in blackjack and NOT ahead in roulette OR NOT ahead in blackjack and ahead in roulette OR ahead in blackjack and ahead in roulette)
= P(ahead in blackjack and NOT ahead in roulette) + P(NOT ahead in blackjack and ahead in roulette) + P(ahead in blackjack and ahead in roulette)
= (0.3 x 0.6) + (0.7 x 0.4) + (0.3 x 0.4)
= 0.18 + 0.28 + 0.12
= 0.58

Cheers,
Brent

RELATED VIDEOS
- The Complement: https://www.gmatprepnow.com/module/gmat ... /video/744
- Probability of Event A OR Event B: https://www.gmatprepnow.com/module/gmat ... /video/747
- Probability of Event A AND Event B: https://www.gmatprepnow.com/module/gmat ... /video/750
- Rewriting Questions: https://www.gmatprepnow.com/module/gmat ... /video/754
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion